40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Aug 6 19:53:15 2013, Jenne, Update, LSC, Arms locked in IR, aligned. IFO at nominal power 
    Reply  Wed Aug 7 10:09:30 2013, Koji, Update, LSC, Arms locked in IR, aligned. IFO at nominal power 
       Reply  Wed Aug 7 21:52:11 2013, Jenne, Update, SUS, SRM coils fine - problem with slow bias slider 
          Reply  Fri Aug 9 21:05:28 2013, Koji, Update, SUS, fixed: SRM coils fine - problem with slow bias slider 
Message ID: 8974     Entry time: Tue Aug 6 19:53:15 2013     Reply to this: 8975
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Arms locked in IR, aligned. IFO at nominal power 

[Koji, Manasa, Jenne]

The Y arm was locked in IR, and we saw flashing in the Xarm (Gautam had the Xarm for green work when we began).  I checked IPANG, and the beam was beautifully unclipped, almost perfectly centered on the first out of vacuum mirror.  I aligned the beam onto the QPD.

We then swapped out the MC Y1 that we use at low power, and replace the usual 10% BS, so that we wouldn't crispy-fry MC REFL.  Manasa adjusted the half wave plate after the laser, to maximize the power going toward the PMC.  We relocked the PMC, and see transmission of ~0.84, which is at the high side of what we usually get.  The beam was aligned onto MC REFL and centered on the WFS, and the MC was locked at nominal power.  Koji tweaked up the alignment of the MC, and ran the WFS offset script.  I aligned beam onto POP QPD and POP110 coarsely (using a flashing PRC, not a locked PRM-ITMY cavity, so the alignment should be rechecked).  The arms have both been locked and aligned in IR....the green beams need to be steered to match the current cavity axis. 

The AS beam, as well as REFL and POP, are all coming out of the vacuum nicely unclipped. 

Notes:  When Koji was aligning the SRM to get the SRC cavity roughly aligned (the AS flashes all overlapping), we noticed that there is some major pitch-yaw coupling.  Serious enough that we should be concerned that perhaps some connector is loose, or an actuator isn't working properly.  This should be checked.

Moral of the story:  Coarse alignment of all mirrors is complete after pump-down and we have IR locked and aligned to both arms at nominal power.

 

Still to do:

* Restore PRM, align beam onto the REFL PDs. 

* Lock PRM-ITMY cavity, align beam onto POP PDs.

* Align AS beam onto AS55. 

* Recenter all oplevs.

* Recenter IPPOS and IPPANG at nominal power.

* Start locking!!

ELOG V3.1.3-